# NPHardEval4V: A Dynamic Reasoning Benchmark of Multimodal Large Language Models

### v2

@@
Abstract
Understanding the reasoning capabilities of Multimodal Large Language Models (MLLMs) is an important area of research. In this study, we introduce a dynamic benchmark, NPHardEval4V, aimed at addressing the existing gaps in evaluating the pure reasoning abilities of MLLMs. Our benchmark aims to provide a venue to disentangle the effect of various factors such as image recognition and instruction following, from the overall performance of the models, allowing us to focus solely on evaluating their reasoning abilities. It is built by converting textual description of questions from NPHardEval to image representations. Our findings reveal significant discrepancies in reasoning abilities across different models and highlight the relatively weak performance of MLLMs compared to LLMs in terms of reasoning. We also investigate the impact of different prompting styles, including visual, text, and combined visual and text prompts, on the reasoning abilities of MLLMs, demonstrating the different impacts of multimodal inputs in model performance. Unlike traditional benchmarks, which focus primarily on static evaluations, our benchmark will be updated monthly to prevent overfitting and ensure a more authentic and fine-grained evaluation of the models. We believe that this benchmark can aid in understanding and guide the further development of reasoning abilities in MLLMs. The benchmark dataset and code are available at https://github.com/lizhouf/NPHardEval4V.
Figure 1: Multimodal Large Language Models’s performance on recognition (RA), Instruction-following (ER), and reasoning (AA) on polynomial time, NP-complete, and NP-hard problems.
1 Introduction
@@
The evolution of Multimodal Large Language Models (MLLMs) marks a significant milestone in the pursuit of artificial general intelligence (AGI), following the advancement of Large Language Models (LLMs) fan2023bibliometric ; zhao2023survey . It introduces new capabilities for understanding and generating content that spans both text and visual inputs, contributing to enhanced multimedia interaction systems and sophisticated cross-modal decision-making tools yin2023survey ; wang2024exploring . Reasoning is a critical ability for MLLMs as it is a fundamental aspect of problem-solving and task completion. The ability to reason enables MLLMs to understand complex relationships between different modalities, draw logical conclusions, and make informed decisions based on the information available.
The assessment of reasoning abilities in MLLMs is a crucial aspect of evaluating their overall performance and guiding the development of more advanced models. For instance, MLLMs can be tasked with reasoning about the minimum number of operations required to transform one string into another by analyzing the visual representation of the two strings and identifying the differences between them. Similarly, given a map, MLLMs can be tasked about reasoning on the optimal route between two locations on a map by analyzing the visual representation of the map and identifying the shortest path based on the available information. The ability to reason enables MLLMs to draw logical conclusions and make informed decisions based on the information available. In this paper, we introduce a benchmark that focuses solely on evaluating the reasoning abilities of various MLLMs. By assessing the extent of their reasoning capabilities, we aim to provide valuable insights into the strengths and limitations of current MLLMs and guide future research towards developing more advanced models with enhanced reasoning abilities. Our benchmark aims to disentangle the effect of various factors such as image recognition and instruction following, from the overall performance of the models, allowing us to focus solely on evaluating their reasoning abilities.
Numerous benchmarks have been developed to assess the capabilities of MLLMs in various domains, such as visual question answering antol2015vqa ; goyal2017making ; marino2019ok ; hudson2019gqa , Optical Character Recognition liu2023hidden , robustness zhao2024evaluating , hallucination liu2023hallusionbench , and holistic overall performance xu2023lvlm ; li2023seed ; liu2023mmbench ; yue2023mmmu . However, despite the breadth of these evaluations, none of them specifically focus on assessing the pure reasoning abilities of MLLMs, leaving a significant gap in our understanding of their reasoning capabilities. Furthermore, many of these benchmarks are static in nature, making them prone to overfitting and limiting their effectiveness in providing a comprehensive view of MLLMs’ abilities chang2023survey . To address these limitations, there is a need for a dynamic benchmark that specifically targets the evaluation of MLLMs’ reasoning abilities and updates regularly to prevent overfitting.
In response to these limitations, we introduce NPHardEval4V, a dynamic benchmark specifically designed to quantitatively and rigorously evaluate the reasoning ability of MLLMs across a diverse set of tasks. This benchmark aims to provide a rigorous framework for assessing MLLMs’ reasoning, leveraging the computational complexity hierarchy to explore the depths of models’ reasoning abilities. The dynamic updating mechanism also makes the benchmark robust in a way that static benchmarks cannot. Moreover, the dynamic nature of NPHardEval4V — with datasets that evolve over time — mitigates the risk of overfitting, ensuring that assessments remain relevant and challenging. This is crucial for fostering models that are genuinely capable of learning and adapting, rather than merely optimizing for static benchmarks.
@@
The proposed benchmark is developed based on the NPHardEval benchmark, as presented in fan2023nphardeval . The NPHardEval benchmark comprises nine types of algorithmic problems, categorized into three polynomial time problems, three NP-complete problems, and three NP-hard problems. Each algorithmic problem consists of 100 instances with varying difficulty levels. To enable a direct comparison between MLLMs and LLMs, we have retained the problems from NPHardEval and converted their textual descriptions into visual representations.
Our research questions are as follow:
  1. 1.
    Reasoning Performance Evaluation of MLLMs: This study seeks to understand the variation in reasoning abilities among different MLLMs and to identify the factors that contribute to any observed performance gaps. By analyzing and separating the influence of recognition and instruction following, we evaluate the core reasoning abilities of MLLMs vary with the nature of complexity and difficulty of problems. This investigation will provide insights into the relative strengths and weaknesses of MLLMs in reasoning tasks. We provide the model details of the MLLMs we evaluate in Table 2.
  2. 2.
    Effects of Vision Input on MLLMs’ Performance: We investigate how the inclusion of vision prompts affects MLLMs’ reasoning abilities compared to pure text inputs. Through ablation studies of three prompt types, including “Figures with Limited Instructional Text Setup" (figure+limited_text), “Text-only Setup" (full_text_only), and “Vision-rich-text Setup" (figure+full_text), this research will assess the impact of vision inputs on models’ reasoning performance and examine how the combination of vision and text inputs can affect MLLMs’ performance.
2 Related Work
@@
2.1 Multimodal Large Language Models (MLLMs) and Their Reasoning Abilities
@@
MLLMs wang2024exploring can process and interpret various multimodal data streams, including imagery and textual content wu2023multimodal . As such, MLLMs are able to surpass singularly moded LLMs and unlock new avenues for performing real-world applications. As summarized by Yin et al. (2023) yin2023survey , MLLMs are more intelligent, user-centric, and holistic as compared to their LLM counterparts. MLLMs can mimic the way of how humans perceive the environment by assimilating multi-sensory inputs that can complement each other. They can also foster intuitive user interactions and communications. The scope of tasks that MLLMs can assist with is significantly broader compared to LLMs, reinforcing their versatility in applications yin2023survey such as engineering cui2024survey and healthcare pal2024gemini .
@@
Reasoning is one of the fundamental intelligent behaviors, essential for solving complex real-world tasks yu2023nature ; huang2022towards . However, even in text-centric settings, LLMs lack proper reasoning abilities, such as dealing with NP-hard (nondeterministic polynomial time) mathematical problems fan2023nphardeval . The quest to explore and improve reasoning capabilities of Strong AI particularly within MLLMs remains a persistent challenge and a pursuit of ongoing research chen2023llava ; liu2023controlllm ; gong2023mindagent .
@@
In the domain of MLLMs, researchers have explored a range of techniques, such as instruction-tuning and prompt engineering, to enhance multimodal reasoning wang2024exploring . The practice of instruction tuning, vital for in-context learning (ICL), has emerged as one of the key techniques in enhancing the reasoning abilities of these models. For example, frozen LLM is an initial MLLM to showcase ICL ability tsimpoukelli2021multimodal . Later, the Flamingo model demonstrated strong ICL capabilities with a more sophisticated LLM coupled with massive-scale image and text data for its pre-training phase alayrac2022flamingo . Prompt engineering, on the other hand, has illustrated its effectiveness in improving the reasoning abilities of MLLMs koh2023grounding ; guo2023images . As summarized by Wang et al. (2024) wang2024exploring , this approach entails a variety of strategic implementations, such as representation learning koh2023grounding , exemplar generation guo2023images , and model interactions wang2022language .
2.2 Benchmarks of Multimodal Large Language Models (MLLMs)
@@
As the reasoning ability of MLLMs continues to advance, benchmarks have become instrumental in evaluating their performance and identifying areas that require improvement. Wang et al. (2024) suggest that a robust multimodal reasoning benchmark must fulfill three key criteria: (1) the integration of multimodal information, (2) the categorization of reasoning, and (3) in-depth annotations of the reasoning steps wang2024exploring .
@@
Previous research has established numerous benchmarks to gauge the performance of MLLMs fu2023mme ; fu2023challenger ; li2023seed ; liu2023hallusionbench ; cao2023towards . Among these, Fu et al. (2023) introduced an extensive MLLM evaluation benchmark to assess models’ perception and cognition skills across 14 distinct subtasks, including commonsense reasoning and code reasoning fu2023mme . Later, Li et al. (2023) presented a benchmark called SEED-Bench, comprising 19,000 multiple-choice questions with precise human annotations. SEED-Bench evaluates MLLMs across 12 dimensions, capturing the ability to understand image and video modalities li2023seed . In a more recent study, Zhang et al. (2024) examined the self-consistency of MLLM responses in the presence of common corruptions. They created MMCBench, an extensive benchmark encompassing over 100 prominent LLMs. This benchmark assesses cross-modal interactions among text, images, and speech, incorporating four key generative tasks: text-to-image, image-to-text, text-to-speech, and speech-to-text zhang2024benchmarking .
@@
Although existing benchmarks have evaluated the abilities of MLLMs across various dimensions, they do not provide a clear picture of the pure reasoning ability of MLLMs because factors such as recognition, knowledge amount, instruction following, and others are all combined in the presented performances in these benchmarks. Given that MLLMs’ general performance are intricately dependent on their recognition process and instruction-following ability, our study aims to factor out other factors in order to see the pure reasoning process and assess them of MLLMs. Moreover, these benchmarks lack dynamic updating mechanisms, which increases the risk of MLLMs becoming overfitted to these benchmarks and restricts their ability to accurately reflect the full range of MLLMs’ capabilities chang2023survey . To address this research gap, our study proposes a dynamic assessment framework for evaluating the reasoning performance of MLLMs. In addition, we utilize the computational complexity hierarchy to rigorously assess the extent to which these MLLMs can achieve in reasoning tasks fan2023nphardeval .
3 Benchmark Construction
@@
As we mentioned above, NPHardEval4V is built upon NPHardEval fan2023nphardeval , using the same set of problems while transformed the input type from textual to visual. This section outlines the transformation of the NPHardEval benchmark to suit the evaluation of MLLMs within NPHardEval4V, emphasizing the multimodal aspects of the tasks and the dynamic nature of the challenges.
3.1 NPHardEval Benchmark and Our Transformation
@@
We explore and build upon the NPHardEval Benchmark fan2023nphardeval , a framework that segments task complexity into three primary computational complexity classes: P (polynomial time), NP-complete (nondeterministic polynomial-time complete), and NP-hard. These classifications serve to delineate the intrinsic difficulty and the computational resources required for solving tasks within each class, showcasing an increasing order of complexity. NPHardEval4V also adopts the same three computational complexity levels (P, NP-Complete, and NP-Hard), transforming textual description of the problems into visual representations [^1] Robustness experiment on the benchmark is performed in fan2023nphardeval ..
@@
To provide a comprehensive overview of task complexity and difficulty within these classes, we introduce a hierarchical categorization illustrated in Table 1, where tasks are further divided into 10 progressive difficulty levels. This granular difficulty grading enables a nuanced assessment of model performance across a spectrum of computational challenges, thereby offering valuable insights into the capabilities and limitations of current models’ reasoning abilities.
@@
Sorted Array Search (SAS)
Table 1: Complexity classes and tasks
To enable the direct comparison with the text-only prompt input in the NPHardEval benchmark, we transform the text-only questions and represent the data part with figures. We further illustrate these transformations, with the data of the problem presented both textually and visually.
Graph Data Transformation
The general construction of graph data problems involves providing MLLMs with both a textual description and a visual representation. For example, in the Graph Coloring Problem (see Figure 2), MLLMs are given a textual prompt alongside a figure depicting a graph, challenging them to reason across modalities. The figure is generated using Python based on the textual description of the graph coloring problem (GCP): it includes information about the nodes and the connections between them.

### v3

# NPHardEval4V: Dynamic Evaluation of Large Vision-Language Models with Effects of Vision

@@
Abstract.
Large Vision-Language Models (LVLMs) have demonstrated impressive capabilities in multimodal understanding, yet their reasoning abilities remain underexplored. Existing benchmarks tend to focus on perception or text-based comprehension, offering limited insight into how well these models perform on structured, logic-driven tasks that require both visual and linguistic reasoning. To address this gap, we introduce NPHardEval4V, a multimodal benchmark suite grounded in four classical NP-hard problems: Knapsack, Set Cover, Traveling Salesperson, and Vertex Cover. Each task is presented through a combination of structured visual layouts and textual prompts, designed to assess the ability of LVLMs to perform combinatorial reasoning under visual-linguistic constraints. We evaluate a set of advanced open-source and closed-source vision-language models under a unified prompting and problem representation framework. This enables fair comparison across models and task types, while isolating key variables affecting performance. Our results show that while these models perform reasonably well on perception-based inputs, they struggle with global optimization, abstraction, and constraint satisfaction. No single model demonstrates consistent reasoning capability across all problem types, and common failure patterns reveal fundamental limitations in current architectures. By leveraging the structure and complexity of NP-hard problems, NPHardEval4V provides a scalable, interpretable, and challenging testbed for diagnosing reasoning behaviors in LVLMs. We hope this benchmark can support the community in building more robust, inference-capable multimodal systems. The benchmark dataset and code are available at https://github.com/lizhouf/NPHardEval4V.
large language model, large vision model, multimodal foundation model, benchmark, complexity classes
journal: TISTcopyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Computing methodologies Information extraction
1. Introduction
@@
The evolution of Large Vision-Language Models (LVLMs) represents a significant milestone in the pursuit of artificial general intelligence (AGI), closely following advancements in Large Language Models (LLMs) (Fan et al., 2024b; Zhao et al., 2023). LVLMs introduce advanced capabilities for understanding and generating integrated textual and visual content, substantially enhancing multimedia interaction systems and cross-modal decision-making applications (Yin et al., 2023; Wang et al., 2024b). Reasoning, a fundamental capability required to interpret complex relationships across modalities, draw logical conclusions, and make informed decisions, is particularly critical for these models.
Evaluating the reasoning abilities of LVLMs is essential yet presents notable challenges. Reasoning tasks demand models to discern intricate relationships between visual and textual information. For instance, models may be tasked with analyzing visual differences between strings to determine transformation operations or interpreting maps to select optimal routes. Numerous existing benchmarks have evaluated LVLMs across various tasks, such as visual question answering (Antol et al., 2015; Goyal et al., 2017; Marino et al., 2019; Hudson and Manning, 2019), Optical Character Recognition (Liu et al., 2023d), robustness (Zhao et al., 2024), hallucination detection (Liu et al., 2023b), and general performance evaluation (Xu et al., 2023; Li et al., 2023; Liu et al., 2023a; Yue et al., 2023). However, these benchmarks have not specifically targeted pure reasoning capabilities, leaving a crucial evaluation gap. Existing tests often lack the rigorous, multi-step logical structure required to truly probe the limits of a model’s reasoning. Additionally, static benchmarks are vulnerable to overfitting, limiting their ability to comprehensively evaluate the evolving capabilities of the models (Chang et al., 2023).
This paper introduces NPHardEval4V, a dynamic benchmark explicitly designed to rigorously evaluate the reasoning capabilities of LVLMs across diverse tasks. NP-hard problems, due to their rigorous mathematical structure, complex constraints, and often non-unique solutions, are presented as naturally suitable and stringent evaluation criteria for visual-language model reasoning. Their inherent difficulty and combinatorial nature make them superior to existing multimodal VQA tests for assessing deep, logical inference. Crucially, these problems are highly representative of real-world applications that require complex optimization and decision-making, such as logistics (Traveling Salesperson Problem), resource allocation (Knapsack Problem), and scheduling (Graph Coloring). Grounding the benchmark in these tasks allows for a more effective measurement of the practical reasoning capabilities of LVLMs. The benchmark’s dynamic nature further ensures a robust evaluation by mitigating overfitting risks.
@@
The proposed benchmark is developed based on the NPHardEval benchmark, as presented in (Fan et al., 2024a). The NPHardEval benchmark comprises nine types of algorithmic problems, categorized into three polynomial time problems, three NP-complete problems, and three NP-hard problems. Each algorithmic problem consists of 100 instances with varying difficulty levels. To enable a direct comparison between LVLMs and LLMs, we have retained the problems from NPHardEval and converted their textual descriptions into visual representations.
We evaluate a set of state-of-the-art LVLMs, including both open- and closed-source models, across multiple reasoning tasks derived from these algorithmic problems. Our evaluation framework incorporates a structured analysis of model performance using metrics such as Recognition Accuracy (RA), Instruction-following Effective Rate (ER), and Aggregated Accuracy (AA). Furthermore, we analyze how reasoning ability varies across different complexity classes and problem difficulty levels, offering a comprehensive view into the capabilities and limitations of existing LVLMs.
Our analysis also reveals consistent patterns in model performance based on input modality. While some models benefit from concise visual prompts, others demonstrate improved reasoning when provided with richer textual context. Notably, we observe that a small subset of models, such as Gemini, exhibit robust performance across multiple input configurations, indicating advanced multimodal integration capabilities. These findings provide important insights into prompt design and model alignment for future research.
The main contributions of this work are as follows:
  1. (1)
    Reasoning Performance Evaluation of LVLMs: We systematically evaluate the reasoning performance of various Large Vision-Language Models (LVLMs), isolating reasoning ability from confounding factors such as image recognition and instruction-following. This analysis reveals the relative strengths and weaknesses across different models. We provide the model details of the LVLMs we evaluate in Table 2.
  2. (2)
    Effects of Vision Input on LVLMs’ Performance: We conduct a comprehensive ablation study on the effect of input modality, comparing three prompt types: (1) visual figures with limited textual instruction (figure+limited_text), (2) text-only descriptions (full_text_only), and (3) vision-rich textual combinations (figure+full_text). This evaluation highlights the impact of visual input on LVLMs’ reasoning performance.
The rest of this paper is organized as follows. Section 2 surveys related work in multimodal reasoning evaluation and benchmark design. Section 3 introduces the construction and multimodal transformation of the proposed benchmark, NPHardEval4V. In Section 4, we describe the experimental settings, including model selection, input types, and evaluation metrics. Section 5 presents experimental results and comparative analyses, followed by an examination of the influence of different input formats in Section 6. Finally, Section 7 discusses key findings, limitations, and future research directions.
2. Related Work
Evaluating multimodal reasoning needs benchmarks beyond perception or association. Although LVLMs perform well on captioning and VQA, they often rely on shallow heuristics. Current benchmarks rarely test structured, symbolic, or constraint-based reasoning that reflects real cognition. This section reviews related benchmarks and LVLM evaluations on complex tasks.
@@ 2. Related Work
2.1. Large Vision-Language Models (LVLMs) and Their Reasoning Abilities
@@
LVLMs (Wang et al., 2024b) can process and interpret various multimodal data streams, including imagery and textual content (Wu et al., 2023). As such, LVLMs are able to surpass singularly moded LLMs and unlock new avenues for performing real-world applications. LVLMs are also more intelligent, user-centric, and holistic as compared to their LLM counterparts (Yin et al., 2023). LVLMs can mimic the way of how humans perceive the environment by assimilating multi-sensory inputs that can complement each other. They can also foster intuitive user interactions and communications. The scope of tasks that LVLMs can assist with is significantly broader compared to LLMs, reinforcing their versatility in applications (Yin et al., 2023) such as engineering (Cui et al., 2024) and healthcare (Pal and Sankarasubbu, 2024).
@@
Reasoning is one of the fundamental intelligent behaviors, essential for solving complex real-world tasks (Yu et al., 2023; Huang and Chang, 2022). However, even in text-centric settings, LLMs lack proper reasoning abilities, such as dealing with NP-hard (nondeterministic polynomial time) mathematical problems (Fan et al., 2024a). The quest to explore and improve reasoning capabilities of Strong AI particularly within LVLMs remains a persistent challenge and a pursuit of ongoing research (Chen et al., 2023; Liu et al., 2023c; Gong et al., 2023).
@@
In the domain of LVLMs, researchers have explored a range of techniques, such as instruction-tuning and prompt engineering, to enhance multimodal reasoning (Wang et al., 2024b). The practice of instruction tuning, vital for in-context learning (ICL), has emerged as one of the key techniques in enhancing the reasoning abilities of these models. For example, frozen LLM is an initial MLLM to showcase ICL ability (Tsimpoukelli et al., 2021). Later, the Flamingo model demonstrated strong ICL capabilities with a more sophisticated LLM coupled with massive-scale image and text data for its pre-training phase (Alayrac et al., 2022). Prompt engineering, on the other hand, has illustrated its effectiveness in improving the reasoning abilities of LVLMs (Koh et al., 2023; Guo et al., 2023). As summarized by (Wang et al., 2024b), this approach entails a variety of strategic implementations, such as representation learning (Koh et al., 2023), exemplar generation (Guo et al., 2023), and model interactions (Wang et al., 2022).
2.2. Benchmarks of Large Vision-Language Models (LVLMs)
@@
As the reasoning ability of LVLMs continues to advance, benchmarks have become instrumental in evaluating their performance and identifying areas that require improvement. A robust multimodal reasoning benchmark must fulfill three key criteria: (1) the integration of multimodal information, (2) the categorization of reasoning, and (3) in-depth annotations of the reasoning steps (Wang et al., 2024b).
@@
Previous research has established numerous benchmarks to gauge the performance of LVLMs (Fu et al., 2023a, b; Li et al., 2023; Liu et al., 2023b; Cao et al., 2023). Among these, (Fu et al., 2023a) introduced an extensive MLLM evaluation benchmark to assess models’ perception and cognition skills across 14 distinct subtasks, including commonsense reasoning and code reasoning. Later, (Li et al., 2023) presented a benchmark called SEED-Bench, comprising 19,000 multiple-choice questions with precise human annotations. SEED-Bench evaluates LVLMs across 12 dimensions, capturing the ability to understand image and video modalities. In a more recent study, (Zhang et al., 2024) further examined the self-consistency of MLLM responses in the presence of common corruptions. They created MMCBench, an extensive benchmark encompassing over 100 prominent LLMs. This benchmark assesses cross-modal interactions among text, images, and speech, incorporating four key generative tasks: text-to-image, image-to-text, text-to-speech, and speech-to-text.
@@
Although existing benchmarks have evaluated the abilities of LVLMs across various dimensions, they do not provide a clear picture of the pure reasoning ability of LVLMs because factors such as recognition, knowledge amount, instruction following, and others are all combined in the presented performances in these benchmarks. Given that LVLMs’ general performance are intricately dependent on their recognition process and instruction-following ability, our study aims to factor out other factors in order to see the pure reasoning process and assess them of LVLMs. Moreover, these benchmarks lack dynamic updating mechanisms, which increases the risk of LVLMs becoming overfitted to these benchmarks and restricts their ability to accurately reflect the full range of LVLMs’ capabilities (Chang et al., 2023). To address this research gap, our study proposes a dynamic assessment framework for evaluating the reasoning performance of LVLMs. In addition, we utilize the computational complexity hierarchy to rigorously assess the extent to which these LVLMs can achieve in reasoning tasks (Fan et al., 2024a).
3. NPHardEval4V: Benchmark Design
@@
NPHardEval4V Benchmark is built on the NPHardEval Benchmark (Fan et al., 2024a), using the same set of problems while transforming the input type from textual to visual. This section first introduced the original NPHardEval Benchmark, including its question hierarchy and rotustness, and then outlines the transformation of the NPHardEval benchmark to suit the evaluation of LVLMs within NPHardEval4V, emphasizing the multimodal aspects of the tasks and the dynamic nature of the challenges.
3.1. NPHardEval Benchmark
@@
The NPHardEval Benchmark segments reasoning task into three primary computational complexity classes: P (polynomial time), NP-complete (nondeterministic polynomial-time complete), and NP-hard. These classifications serve to delineate the intrinsic difficulty and the computational resources required for solving tasks within each class, showcasing an increasing order of complexity.
@@
To provide a comprehensive overview of task complexity and difficulty within these classes, we introduce a hierarchical categorization illustrated in Table 1, where tasks are divided into 10 levels of progressive difficulty. This granular difficulty grading enables a nuanced assessment of model performance across a spectrum of computational challenges, thereby offering valuable insights into the capabilities and limitations of current models’ reasoning abilities.
Table 1. Complexity classes and tasks
@@
Binary Search Problem (BSP)
For each level, we then randomly generate 10 questions and this results in a total of 900 questions for one version in the dynamic benchmark. The benchmark is designed to be comprehensive and is supported by a dynamic update mechanism that ensures continuous renewal when necessary. While an infinite number of reasoning questions would ideally be required to precisely evaluate a model’s capability on a specific task, practical constraints necessitate a finite sample. Based on sample size estimations (MacCallum et al., 1999; Morse, 1992) and margin of error calculations (Israel et al., 1992; Yamane, 1973), 100 questions per task are necessary to achieve a 95% confidence level with a 10% margin of error.
3.2. Transformation to NPHardEval4V
To enable direct comparison with the text-only prompt input in the NPHardEval benchmark, we transform the text-only questions and represent the data part with figures. We further illustrate these transformations, with the data of the problem presented both textually and visually.
3.2.1. Graph Data Transformation
The general construction of graph data problems involves providing LVLMs with both a textual description and a visual representation. For example, in the Graph Coloring Problem (see Figure 1), LVLMs are given a textual prompt alongside a figure depicting a graph, challenging them to reason across modalities. The figure is generated using Python based on the textual description of the graph coloring problem (GCP): it includes information about the nodes and the connections between them.